OpenStudy (anonymous):

Zaineb is deciding whether to buy a stereo at $695 plus GST and PST (15%) now, or to invest the money to buy the stereo in a year. Her account pays 7.35% per annum, compounded monthly. If she can buy the stereo at the same price next year, how much would she save by investing the money?

OpenStudy (anonymous):

GST and PST is 15%

OpenStudy (anonymous):

I tried: Zaineb buys a stereo at $695 plus GST and PST now... P = 695 i = 0.15(2) = 0.30 n = 0 A = P(1+i) = 695(1 + 0.30) = 903.5 Zainet invest the money to buy the stereo in a year... P = 695 i = 0.0735/12 = 0.006125 n = 1(12) = 12 A = 695(1 + 0.006125)^12 = 747.84 Interest earned for buying the stereo now = Amount - Principle = 903.5 - 695 = 208.5 Interest earned for buying the stereo next year = Amount - Principle = 747.84 - 695 = 52.84 The answer at the back of my book is $60. 76